Answer problem in full.

Genuine Spice Inc. began operations on January 1 of the current year. The company produces eight-ounce bottles of hand and bo1. Determine the fixed and variable portion of the utility cost using the high-low method. Round the per unit cost to the neaNote: This section is a continuation from Part A of the comprehensive problem. Be sure you have completed Part A before attemPart B-August Budgets During July of the current year, the management of Genuine Spice Inc. asked the controller to prepare A5. Prepare the August production budget. Enter all amounts as positive numbers. Genuine Spice Inc. Production Budget For the8. Prepare the August factory overhead cost budget. If an amount box does not require an entry, leave it blank. Genuine SpiceNote: This section is a continuation from Parts A and B of the comprehensive problem. Be sure you have completed Parts A andPart C-August Variance Analysis During September of the current year, the controller was asked to perform variance analyses f10. Determine and interpret the direct materials price and quantity variances for the three materials. Enter the costs in dolDirect Labor Time Variance: Mixing Filling Department Department Actual time (hours) Standard time (hours) Difference Standar
Answer #1
Genuine Spice Inc.
High Low Method=change in cost/Change In Production
High Low Method= ($740-$600)/(1200-500)
Variable cost per unit= $140/700= $             0.20
1) At Higher level variable cost=(1200*.20) $                                  240.00
Fixed cost=($740-$240) $                                  500.00
Total cost $                                  740.00
At Lower level Variable cost=(500*.20) $                                  100.00
Fixed Cost $                                  500.00
Total cost $                                  600.00
Variable cost per unit will be remain the same at all levels of activity, but fixed cost in total will remain the same.
2) Calculation of contribution Margin per case
Selling Price=(A) $        100.00
Variable cost:
Direct Material($2+$9+$6) $                                    17.00
Direct Labor($6+$1.20) $                                      7.20
Utilities $                                      0.20
Selling Expense-Selling commission $                                    20.00
Total variable cost per case=(B) $          44.40
Contribution Margin per case=(A)-(B) $          55.60
3) Calculation of Fixed cost:
Utilities Expense $                                  500.00
Facility Lease $                            14,000.00
Equipment Depreciation $                              4,300.00
Supplies $                                  660.00
Total $                            19,460.00
4) Break even Sales in Units=(Fixed cost/Contribution Margin per unit)
Fixed Cost=(A) $                            19,460.00
Contribution Margin per unit=(B) $                                    55.60
Break even Sales in Units=(A)/(B) 350 Cases
Part B 5)
Genuine Spice Inc.
Production Budget
For the month ended August 31st,
Expected cases to be sold 1500
Plus: desired ending inventory 175
Cases required 1675
Less: beginning Inventory -300
Cases produced 1375
6) Genuine Spice Inc.
Purchase Budget
For the month ended August 31st,
Cream base Natural Oils Bottles Total
Units required for production(1375*100 ozs),(1375*30 ozs),(1375*12 bottles)=(A) 137500 41250 16500
Add: Ending Inventory=(B) 1000 360 240
Total needs=(C )=(A)+(B) 138500 41610 16740
Less: Beginning Inventory=(D) 250 290 600
Direct Material to be purchases=(E )=(C )-(D) 138250 41320 16140
Unit purchase price=(F) $                                      0.02 $             0.30 $         0.50
Total Direct Material purchased=(E )*(F) $                              2,765.00 $ 12,396.00 $ 8,070.00 $ 23,231.00